OpenStudy (anonymous):

Help with questions 1-5 please?

OpenStudy (anonymous):

1.What is the length of RS? 2. What is the value of x? 3. What is the value of y? 4. What is the length of UW? 5. What is the length of UV?

OpenStudy (amistre64):

there seems to be alot of assumptions wed have to make for that picture to come up with any sort of solutions.

OpenStudy (anonymous):

That's why im so confused

OpenStudy (amistre64):

assume all the triangles are similar to 

OpenStudy (amistre64):

2x = 11 3y = 15.9 and the side with the 12 on it can either be scaled or determine the similar proportion of the top segment

OpenStudy (anonymous):

Okay thank you

OpenStudy (amistre64):

other than that, im as lost at it as you are :) good luck

OpenStudy (anonymous):

Hahaha I'll try my best!

OpenStudy (amistre64):

if im still around, i can dbl chk what you get if you need me to

OpenStudy (anonymous):

Ok thank you, i will post my answers once I'm done

OpenStudy (amistre64):

you say: 1. 24 2. 6 3. 5.3 4. 7.95 5. 12 id say x=11/2, y=16/3 for starters then the proportion\[\frac{12}{11}=\frac{RS}{11+11/2}\] \[\frac{12(11+11/2)}{11}=RS\] \[12(1+1/2)=RS=18\]

OpenStudy (anonymous):

Okayy, i see i see

OpenStudy (amistre64):

but someting just doesnt sit right for me in that .... if 2x=11, then RS should equal 24

OpenStudy (amistre64):

lol, i did 11+11/2 and not 11+2(11/2) :)

OpenStudy (amistre64):

uv = 15.9 uw=11 x=11/2 = 5.5 y=5.3 rs=24 sounds good to me

OpenStudy (anonymous):

Sorry i took so long, went on break but thank you so much :)

OpenStudy (anonymous):

Those were all right :)